### v1

@@ 3 Convergence analysis
To prove Theorem 2.1, we first give several shorthand notation. We denote (r)SVRG iterates for the noisy data $y^{\delta}$ by $x_{k}^{\delta}$. For any $K=0,1,\cdots$ and $t=0,\cdots,M-1$, we define
@@
$\overline{\Phi}_{i}^{i^{\prime}}(j^{\prime},r)$ $=\sum_{j=i}^{i^{\prime}}(i^{\prime}+j^{\prime}-j)^{-r}\mathbb{E}[\|A\Delta_{j}^{\delta}\|^{2}],\quad\Phi_{i}^{i^{\prime}}(j^{\prime},r)=\sum_{j=i}^{i^{\prime}}(i^{\prime}+j^{\prime}-j)^{-r}\|A\Delta_{j}^{\delta}\|,$
$\phi_{i}^{i^{\prime}}$ $=\sum_{j=i}^{i^{\prime}}P^{i^{\prime}-j}N_{j}\Delta_{j}^{\delta},\quad\tilde{\phi}^{i^{\prime}}=\sum_{j=0}^{i^{\prime}}P^{j},\qquad\forall i,\;i^{\prime},\;j^{\prime}\geq 0,$
and follow the conventions $\sum_{j=i}^{i^{\prime}}R_{j}=0$ and $\sum_{j=-i}^{i^{\prime}}R_{j}=\sum_{j=0}^{i^{\prime}}R_{j}$ for any sequence $\{R_{j}\}_{j}$ and $0\leq i^{\prime}<i$, and $0^{s}=1$ for any $s$. Under Assumption 2.1(iii), $A_{\delta}:=A_{\dagger}-A$ satisfies $A^{*}A_{\delta}=0$ and $\|A_{\delta}\|<a\delta^{b}.$ Similarly, let $B_{\dagger}:=\mathbb{E}[A_{{\dagger},i}^{*}A_{{\dagger},i}]=n^{-1}A_{\dagger}^{*}A_{\dagger}$ and $B_{\delta}:=B_{\dagger}-B=n^{-1}A_{\delta}^{*}A_{\delta}$. Then $B^{*}B_{\delta}=0$ and $\|B_{\delta}\|<n^{-1}a^{2}\delta^{2b}.$
@@ 3.1 Error decomposition
For any $K\geq 0$ and $0\leq t\leq M-1$, we decompose the error $e_{KM+t+1}^{\delta}\equiv x_{KM+t+1}^{\delta}-x_{\dagger}$ and the weighted successive error $A\Delta_{KM+t}^{\delta}=A(x_{KM+t}^{\delta}-x_{KM}^{\delta})$ between the $(KM+t)$th and $KM$th iterations into the bias and variance, which plays a crucial role in the analysis.
@@
Similarly, by Lemma 3.1 and the identity $\mathbb{E}[\langle N_{i},N_{j}\rangle|\mathcal{F}_{i}]=0$ for any $j>i$, we have
@@
Then, by Lemmas A.2 and A.1, we derive
@@
Similarly, by the triangle inequality and Lemmas A.2 and A.1, we obtain
@@ 3.2 Convergence analysis
$\overline{\Phi}_{1}^{k-1}(0,1)\leq 4(c_{1}+c_{2}\delta^{2})k^{-1}.$
Next, we bound the first two terms in (3.3). By the definitions $B_{\dagger}=B+B_{\delta}$ and $P=I-c_{0}B$, and the identity $B^{*}B_{\delta}=0$, Assumption 2.1(ii) implies
@@
Next we bound ${\rm I}:=n^{-\frac{1}{2}}c_{0}\|\tilde{\phi}^{k-1}B^{\frac{1}{2}}\|\delta$. If $a=0$, Lemma A.1 and the triangle inequality imply
${\rm I}\leq$ $n^{-\frac{1}{2}}c_{0}\sum_{j=0}^{k-1}\|P^{j}B^{\frac{1}{2}}\|\delta\leq\sqrt{\frac{c_{0}}{2n}}\sum_{j=0}^{k-1}j^{-\frac{1}{2}}\delta\leq\sqrt{\frac{2c_{0}}{n}}\sqrt{k}\delta;$ (3.5)
@@
This proves the estimate (2.4). Similarly, for $\|e_{k}^{\delta}\|$ when $c_{0}<C_{0}$, Lemma 3.2 yields
@@
The parameter $b$ in Assumption 2.1(iii) is set to $b=(1+2\nu)^{-1}$. Now we discuss the choice of $a$. From the bound on $\|e_{k}^{\delta}\|$ (or $\mathbb{E}[\|e_{k}^{\delta}\|^{2}]^{\frac{1}{2}}$) in the proof of Theorem 2.1, we have
@@
$c_{n,\nu}$ attains its minimum at $a_{*}:=\big(n^{\nu}/(2\nu\|w\|)\big)^{\frac{1}{1+2\nu}}$, and $c_{n,\nu}(a_{*})=\big((2\nu)^{-\frac{2\nu}{1+2\nu}}+(2\nu)^{\frac{1}{1+2\nu}}\big)n^{-\frac{\nu}{1+2\nu}}\|w\|^{\frac{1}{1+2\nu}}$. To avoid the blow-up of $a$ as $\nu\to 0^{+}$, let $a=(n^{\nu}/\|w\|)^{\frac{1}{1+2\nu}}$. Then $c_{n,\nu}(a)=2n^{-\frac{\nu}{1+2\nu}}\|w\|^{\frac{1}{1+2\nu}}$.
Next we prove Corollary 2.2, which relax Assumption 2.1(iii).
@@
$\|P^{k}e_{0}^{\delta}\|=$ $\|P^{k}B_{\dagger}^{\nu}w\|\leq\|P^{k}B^{\nu}w\|+\|P^{k}(B_{\dagger}^{\nu}-B^{\nu})w\|\leq\|P^{k}B^{\nu}w\|+\|B_{\dagger}^{\nu}-B^{\nu}\|\|w\|.$
When $\nu\in(0,1]$, by [17, Theorem 2.3], the term ${\rm I}:=\|B_{\dagger}^{\nu}-B^{\nu}\|$ can be bounded by
${\rm I}\leq$ $\|B_{\dagger}-B\|^{\nu}=n^{-\nu}\|A_{\dagger}^{*}A_{\dagger}-A^{*}A\|^{\nu}\leq n^{-\nu}\big(\|A_{\dagger}^{*}\|\|A_{\dagger}-A\|+\|A_{\dagger}^{*}-A^{*}\|\|A\|\big)^{\nu}$
@@
$\leq$ $n^{-(\nu-1)}\nu\big(\|A_{\dagger}\|+\epsilon_{A}\big)^{2(\nu-1)}\|B_{\dagger}-B\|\leq 2^{2\nu}n^{-\nu}\nu\|A_{\dagger}\|^{2\nu-1}\epsilon_{A}.$
Then, let $\epsilon_{A}\leq\delta^{\frac{2\nu}{(1+2\nu)\min(1,\nu)}}$, we have ${\rm I}\leq c(\nu)n^{-\nu}\epsilon_{A}\leq c(\nu)n^{-\nu}\delta^{\frac{2\nu}{1+2\nu}}$, with the constant $c(\nu)$ independent of $\delta$ and $n$. The assumption on $A$ implies $\|A^{-1}\|\leq a^{-1}\delta^{-b}$ or the nonzero singular values $\sigma$ of $A$ such that $\sigma\geq a\delta^{b}>0$, which implies (3.6). Thus, Theorem 2.1 still holds. ∎
@@
If $A_{\dagger}$ is compact, with its SVD $A_{\dagger}(\cdot)=\sum_{j=1}^{\infty}\sigma_{j}\langle\varphi_{j},\cdot\rangle\psi_{j}$, where the singular values $\{\sigma_{j}\}_{j=1}^{\infty}$ such that $\sigma_{j}\geq\sigma_{j^{\prime}}\geq 2a\delta^{b}>0$ for any $j\leq j^{\prime}\leq J$ and $\sigma_{j}<2a\delta^{b}$ for any $j>J$. For any small $\epsilon_{A}\in(0,a\delta^{b})\subset(0,\|A^{\dagger}\|)$, we may approximate $A_{\dagger}$ by $A(\cdot)=\sum_{j=1}^{J}\tilde{\sigma}_{j}\langle\tilde{\varphi}_{j},\cdot\rangle\tilde{\psi}_{j}$ with $\{\tilde{\varphi}_{j}\}_{j=1}^{J}$ and $\{\tilde{\psi}_{j}\}_{j=1}^{J}$ being orthonormal in $X$ and $Y$, respectively, which satisfies $\|\varphi_{j}-\tilde{\varphi}_{j}\|<\epsilon_{A}$ and $|\tilde{\sigma}_{j}-\sigma_{j}|<\epsilon_{A}$. Then we take $A_{\rm T}(\cdot)=\sum_{j=1}^{J}\sigma_{j}\langle\varphi_{j},\cdot\rangle\psi_{j}$. Let $B(\cdot)=n^{-1}A^{*}A(\cdot)=\frac{1}{n}\sum_{j=1}^{J}\tilde{\sigma}_{j}^{2}\langle\tilde{\varphi}_{j},\cdot\rangle\tilde{\varphi}_{j}$, $B_{\rm T}(\cdot)=n^{-1}A_{\rm T}^{*}A_{\rm T}(\cdot)=n^{-1}\sum_{j=1}^{J}\sigma_{j}^{2}\langle\varphi_{j},\cdot\rangle\varphi_{j}$, and $B_{\rm T,\delta}=B_{\dagger}-B_{\rm T}$. Then there hold $B_{\rm T}^{*}B_{\rm T,\delta}=0$ and $\|B_{\rm T,\delta}\|<4n^{-1}a^{2}\delta^{2b}$. Hence,
${\rm I}$ $=\|B_{\dagger}^{\nu}-B^{\nu}\|\leq\|(B_{\rm T}+B_{\rm T,\delta})^{\nu}-B^{\nu}\|=\|B_{\rm T}^{\nu}+B_{\rm T,\delta}^{\nu}-B^{\nu}\|\leq\|B_{\rm T,\delta}\|^{\nu}+{\rm I}_{\epsilon},$
with ${\rm I}_{\epsilon}:=n^{-\nu}\sup_{\|z\|=1}\big\|\sum_{j=1}^{J}\big(\sigma_{j}^{2\nu}\langle\varphi_{j},z\rangle\varphi_{j}-\tilde{\sigma}_{j}^{2\nu}\langle\tilde{\varphi}_{j},z\rangle\tilde{\varphi}_{j}\big)\big\|$. By the triangle inequality,
${\rm I}_{\epsilon}\leq$ $n^{-\nu}\sup_{\|z\|=1}\bigg\|\sum_{j=1}^{J}\sigma_{j}^{2\nu}\langle\varphi_{j},z\rangle(\varphi_{j}-\tilde{\varphi}_{j})\bigg\|+n^{-\nu}\sup_{\|z\|=1}\bigg\|\sum_{j=1}^{J}(\sigma_{j}^{2\nu}-\tilde{\sigma}_{j}^{2\nu})\langle\varphi_{j},z\rangle\tilde{\varphi}_{j}\bigg\|$
@@
$\leq$ $n^{-\nu}(\|A_{\dagger}\|^{2\nu}+\|A\|^{2\nu})\epsilon_{A}+n^{-\nu}\sup_{j\leq J}(\sigma_{j}^{2\nu}-\tilde{\sigma}_{j}^{2\nu})$
@@
Note that the initial error $x_{\dagger}-x_{0}\in\overline{\mathrm{Range}(A_{\dagger}^{*})}$. The polar decomposition $A_{\dagger}=Q(A_{\dagger}^{*}A_{\dagger})^{\frac{1}{2}}$ with a partial isometry $Q$ (i.e. $Q^{*}Q$ and $QQ^{*}$ are projections) implies $x_{\dagger}-x_{0}\in\overline{\mathrm{Range}\big((A_{\dagger}^{*}A_{\dagger})^{\frac{1}{2}}\big)}$. Thus, for any $\epsilon_{0}>0$, there exists some $\tilde{x}_{0}$, satisfying Assumption 2.1(ii) with $\nu=\frac{1}{2}$, such that $\|x_{0}-\tilde{x}_{0}\|<\epsilon_{0}$. Let $\tilde{x}^{\delta}_{k}$ be the (r)SVRG iterate starting with $\tilde{x}_{0}$ and $\tilde{e}^{\delta}_{k}=\tilde{x}^{\delta}_{k}-x_{\dagger}$. Then, when $b=\frac{1}{2}$, by Lemma A.1 and the inequality (3.4), we can bound $\|P^{k}e_{0}^{\delta}\|$ in (3.3) and (3.7) by
@@
$\leq\big((2c_{0})^{-\frac{1}{2}}k^{-\frac{1}{2}}+n^{-\frac{1}{2}}a\sqrt{\delta}\big)\|w\|+\epsilon_{0}.$
Consequently,
$\begin{array}[]{cc}\mathbb{E}[\|e_{k}^{\delta}\|^{2}]^{\frac{1}{2}}\leq\epsilon_{0}+c^{*}k^{-\frac{1}{2}}+c^{*}\left\{\begin{array}[]{cc}\sqrt{\delta},&a>0,\\ n^{-\frac{1}{2}}\sqrt{k}\delta,&a=0,\end{array}\right.&c_{0}<\overline{C_{0}}.\\ \|e_{k}^{\delta}\|\leq\epsilon_{0}+\sqrt{n}c^{*}k^{-\frac{1}{2}}\ln k+c^{*}\left\{\begin{array}[]{cc}\sqrt{\delta},&a>0,\\ n^{-\frac{1}{2}}\sqrt{k}\delta,&a=0,\end{array}\right.&c_{0}<C_{0}.\end{array}$
Taking the limit as $\epsilon_{0}\to 0^{+}$ completes the proof of the corollary. ∎

### v2

@@ 3 Convergence analysis
To prove Theorem 2.1, we first give several shorthand notation. We denote the (r)SVRG iterates for the noisy data $y^{\delta}$ by $x_{k}^{\delta}$. For any $K=0,1,\cdots$ and $t=0,\cdots,M-1$, we define
@@
$\phi_{i}^{i^{\prime}}$ $=\sum_{j=i}^{i^{\prime}}P^{i^{\prime}-j}N_{j}\Delta_{j}^{\delta},\quad\tilde{\phi}^{i^{\prime}}=\sum_{j=0}^{i^{\prime}}P^{j},\quad\forall i,\;i^{\prime}\geq 0,$
$\overline{\Phi}_{i}^{i^{\prime}}(0,1)$ $=2c_{0}\|B\|\mathbb{E}[\|A\Delta_{i^{\prime}}^{\delta}\|^{2}]+\sum_{j=i}^{i^{\prime}-1}(i^{\prime}-j)^{-1}\mathbb{E}[\|A\Delta_{j}^{\delta}\|^{2}],\quad\forall i^{\prime}\geq i\geq 0,$
$\Phi_{i}^{i^{\prime}}(0,\tfrac{1}{2})$ $=\sqrt{2c_{0}\|B\|}\|A\Delta_{i^{\prime}}^{\delta}\|+\sum_{j=i}^{i^{\prime}-1}(i^{\prime}-j)^{-\frac{1}{2}}\|A\Delta_{j}^{\delta}\|,\quad\forall i^{\prime}\geq i\geq 0,$
and for any $r,\;j^{\prime}>0$ or $r=0$,
$\overline{\Phi}_{i}^{i^{\prime}}(j^{\prime},r)$ $=\sum_{j=i}^{i^{\prime}}(i^{\prime}+j^{\prime}-j)^{-r}\mathbb{E}[\|A\Delta_{j}^{\delta}\|^{2}],$
$\Phi_{i}^{i^{\prime}}(j^{\prime},r)$ $=\sum_{j=i}^{i^{\prime}}(i^{\prime}+j^{\prime}-j)^{-r}\|A\Delta_{j}^{\delta}\|,$
and follow the conventions $\sum_{j=i}^{i^{\prime}}R_{j}=0$ and $\sum_{j=-i}^{i^{\prime}}R_{j}=\sum_{j=0}^{i^{\prime}}R_{j}$ for any sequence $\{R_{j}\}_{j}$ and $0\leq i^{\prime}<i$. Under Assumption 2.1(iii), $A_{\delta}:=A_{\dagger}-A$ satisfies $A^{*}A_{\delta}=0$ and $\|A_{\delta}\|<a\delta^{b}.$ Similarly, let
$B_{\dagger}:=\mathbb{E}[A_{{\dagger},i}^{*}A_{{\dagger},i}]=n^{-1}A_{\dagger}^{*}A_{\dagger}\quad\mbox{and}\quad B_{\delta}:=B_{\dagger}-B=n^{-1}A_{\delta}^{*}A_{\delta}.$
Then $B^{*}B_{\delta}=0$ and $\|B_{\delta}\|<n^{-1}a^{2}\delta^{2b}.$
@@ 3.1 Error decomposition
For any $K\geq 0$ and $0\leq t\leq M-1$, we decompose the error $e_{KM+t+1}^{\delta}\equiv x_{KM+t+1}^{\delta}-x_{\dagger}$ and the weighted successive error $A\Delta_{KM+t}^{\delta}=A(x_{KM+t}^{\delta}-x_{KM}^{\delta})$ between the $(KM+t)$th and $KM$th iterations into the bias and variance components, which plays a crucial role in the subsequent analysis. (Note that the $KM$th iteration is an anchor point.)
@@
Similarly, by Lemma 3.1 and the identity $\mathbb{E}[\langle P^{k-1-i}N_{i}\Delta_{i}^{\delta},P^{k-1-j}N_{j}\Delta_{j}^{\delta}\rangle|\mathcal{F}_{j}]=0$ for any $j>i$, we have
@@
Then, by Lemmas A.1 and A.2, we derive
@@
Similarly, by the triangle inequality and Lemmas A.1 and A.2, we obtain
@@ 3.2 Convergence analysis
$\overline{\Phi}_{1}^{k-1}(0,1)\leq(3+2c_{0}\|B\|)(c_{1}+c_{2}\delta^{2})k^{-1}.$
Next, we bound the first two terms in (3.3). By the definitions $B_{\dagger}=B+B_{\delta}$, $P=I-c_{0}B$ and $B=n^{-1}A^{*}A$ with $A$ satisfying Assumption 2.1(iii), Assumption 2.1(ii) implies
@@
These identities follow directly from the definition of the fractional power $B_{\dagger}^{\nu}w$ via spectral decomposition:
$B_{\dagger}^{\nu}w=\sum_{j=1}^{\infty}(n^{-1}\sigma_{j}^{2})^{\nu}\langle\varphi_{j},w\rangle\varphi_{j}$
$=$ $\sum_{j=1}^{J}(n^{-1}\sigma_{j}^{2})^{\nu}\langle\varphi_{j},w\rangle\varphi_{j}+\sum_{j=J+1}^{\infty}(n^{-1}\sigma_{j}^{2})^{\nu}\langle\varphi_{j},w\rangle\varphi_{j}=B^{\nu}w+B_{\delta}^{\nu}w.$ (3.4)
@@
Next we bound the term ${\rm I}:=n^{-\frac{1}{2}}c_{0}\|\tilde{\phi}^{k-1}B^{\frac{1}{2}}\|\delta$. If $a=0$, Lemma A.1, the triangle inequality and inequality $c_{0}\|B\|\leq\max(\overline{C_{0}},C_{0})L\leq\frac{1}{2}$ imply
${\rm I}\leq$ $n^{-\frac{1}{2}}c_{0}\sum_{j=0}^{k-1}\|P^{j}B^{\frac{1}{2}}\|\delta\leq\sqrt{\frac{c_{0}}{2n}}\Big(\sqrt{2c_{0}\|B\|}+\sum_{j=1}^{k-1}j^{-\frac{1}{2}}\Big)\delta$
$\leq$ $\sqrt{\frac{c_{0}}{2n}}\Big(1+1+\int_{1}^{k}x^{-\frac{1}{2}}\;{\rm d}x\Big)\delta$
$\leq$ $\sqrt{\frac{c_{0}}{2n}}(2+2\sqrt{k}-2)\delta\leq\sqrt{\frac{2c_{0}}{n}}\sqrt{k}\delta;$ (3.6)
@@
This proves the estimate (2.4). Similarly, for $\|e_{k}^{\delta}\|$, when $c_{0}<C_{0}$, Lemma 3.2 yields
@@
The parameter $b$ in Assumption 2.1(iii) is set to $b=(1+2\nu)^{-1}$. Now we discuss the suitable choice of the parameter $a$. From the bound on $\|e_{k}^{\delta}\|$ (or $\mathbb{E}[\|e_{k}^{\delta}\|^{2}]^{\frac{1}{2}}$) in the proof of Theorem 2.1, we have
@@
Note that $c_{n,\nu}$ attains its minimum at $a_{*}:=\big(n^{\nu}/(2\nu\|w\|)\big)^{\frac{1}{1+2\nu}}$, with
$c_{n,\nu}(a_{*})=\Big((2\nu)^{-\frac{2\nu}{1+2\nu}}+(2\nu)^{\frac{1}{1+2\nu}}\Big)n^{-\frac{\nu}{1+2\nu}}\|w\|^{\frac{1}{1+2\nu}}.$
To avoid the blow-up of $a$ as $\nu\to 0^{+}$, let $a=(n^{\nu}/\|w\|)^{\frac{1}{1+2\nu}}$. Then $c_{n,\nu}(a)=2n^{-\frac{\nu}{1+2\nu}}\|w\|^{\frac{1}{1+2\nu}}$.
Next we prove Corollary 2.2, which relaxes Assumption 2.1(iii).
@@
$\|P^{k}e_{0}^{\delta}\|=$ $\|P^{k}B_{\dagger}^{\nu}w\|\leq\|P^{k}B^{\nu}w\|+\|P^{k}(B_{\dagger}^{\nu}-B^{\nu})w\|$
$\leq$ $\|P^{k}B^{\nu}w\|+\|B_{\dagger}^{\nu}-B^{\nu}\|\|w\|.$
When $\nu\in(0,1]$, by [23, Theorem 2.3], the term ${\rm I}:=\|B_{\dagger}^{\nu}-B^{\nu}\|$ can be bounded by
${\rm I}\leq$ $\|B_{\dagger}-B\|^{\nu}=n^{-\nu}\|A_{\dagger}^{*}A_{\dagger}-A^{*}A\|^{\nu}$
$\leq$ $n^{-\nu}\big(\|A_{\dagger}^{*}\|\|A_{\dagger}-A\|+\|A_{\dagger}^{*}-A^{*}\|\|A\|\big)^{\nu}$
@@
$\leq$ $n^{-(\nu-1)}\nu\big(\|A_{\dagger}\|+\epsilon_{A}\big)^{2(\nu-1)}\|B_{\dagger}-B\|$
$\leq$ $2^{2\nu}\nu n^{-\nu}\|A_{\dagger}\|^{2\nu-1}\epsilon_{A}.$
Then, let $\epsilon_{A}\leq\delta^{\frac{2\nu}{(1+2\nu)\min(1,\nu)}}$, we have
${\rm I}\leq c(\nu)n^{-\nu}\epsilon_{A}\leq c(\nu)n^{-\nu}\delta^{\frac{2\nu}{1+2\nu}},$
with the constant $c(\nu)$ independent of $\delta$ and $n$. The assumption on $A$ implies $\|A^{-1}\|\leq a^{-1}\delta^{-b}$ or the nonzero singular values $\sigma$ of $A$ such that $\sigma\geq a\delta^{b}>0$, which implies (3.7). Thus, Theorem 2.1 still holds. ∎
@@
If $A_{\dagger}$ is compact, with its SVD $A_{\dagger}(\cdot)=\sum_{j=1}^{\infty}\sigma_{j}\langle\varphi_{j},\cdot\rangle\psi_{j}$, where the singular values $\{\sigma_{j}\}_{j=1}^{\infty}$ such that $\sigma_{j}\geq\sigma_{j^{\prime}}\geq 2a\delta^{b}>0$ for any $j\leq j^{\prime}\leq J$ and $\sigma_{j}<2a\delta^{b}$ for any $j>J$. For any small $\epsilon_{A}\in(0,a\delta^{b})\subset(0,\|A^{\dagger}\|)$, we may approximate the operator $A_{\dagger}$ by $A(\cdot)=\sum_{j=1}^{J}\tilde{\sigma}_{j}\langle\tilde{\varphi}_{j},\cdot\rangle\tilde{\psi}_{j}$ with $\{\tilde{\varphi}_{j}\}_{j=1}^{J}$ and $\{\tilde{\psi}_{j}\}_{j=1}^{J}$ being orthonormal in $X$ and $Y$, respectively, which satisfies
$\|\varphi_{j}-\tilde{\varphi}_{j}\|<\epsilon_{A}\quad\mbox{and}\quad|\tilde{\sigma}_{j}-\sigma_{j}|<\epsilon_{A}.$
Then we take $A_{\rm T}(\cdot)=\sum_{j=1}^{J}\sigma_{j}\langle\varphi_{j},\cdot\rangle\psi_{j}$. Let
$B(\cdot)$ $=n^{-1}A^{*}A(\cdot)=n^{-1}\sum_{j=1}^{J}\tilde{\sigma}_{j}^{2}\langle\tilde{\varphi}_{j},\cdot\rangle\tilde{\varphi}_{j},$
$B_{\rm T}(\cdot)$ $=n^{-1}A_{\rm T}^{*}A_{\rm T}(\cdot)=n^{-1}\sum_{j=1}^{J}\sigma_{j}^{2}\langle\varphi_{j},\cdot\rangle\varphi_{j},$
and $B_{\rm T,\delta}=B_{\dagger}-B_{\rm T}$. Then there hold $B_{\rm T}^{*}B_{\rm T,\delta}=0$, $\|B_{\rm T,\delta}\|<4n^{-1}a^{2}\delta^{2b}$, and $B_{\dagger}^{\nu}=B_{\rm T}^{\nu}+B_{\rm T,\delta}^{\nu}$ due to the identity (3.4). Hence,
${\rm I}$ $=\|B_{\dagger}^{\nu}-B^{\nu}\|=\|B_{\rm T}^{\nu}+B_{\rm T,\delta}^{\nu}-B^{\nu}\|\leq\|B_{\rm T,\delta}\|^{\nu}+{\rm I}_{\epsilon},$
with
${\rm I}_{\epsilon}:=n^{-\nu}\sup_{\|z\|=1}\big\|\sum_{j=1}^{J}\big(\sigma_{j}^{2\nu}\langle\varphi_{j},z\rangle\varphi_{j}-\tilde{\sigma}_{j}^{2\nu}\langle\tilde{\varphi}_{j},z\rangle\tilde{\varphi}_{j}\big)\big\|.$
By the triangle inequality, we obtain
${\rm I}_{\epsilon}\leq$ $n^{-\nu}\sup_{\|z\|=1}\bigg\|\sum_{j=1}^{J}\sigma_{j}^{2\nu}\langle\varphi_{j},z\rangle(\varphi_{j}-\tilde{\varphi}_{j})\bigg\|$
$+n^{-\nu}\sup_{\|z\|=1}\bigg\|\sum_{j=1}^{J}(\sigma_{j}^{2\nu}-\tilde{\sigma}_{j}^{2\nu})\langle\varphi_{j},z\rangle\tilde{\varphi}_{j}\bigg\|$
@@
$\leq$ $n^{-\nu}(\|A_{\dagger}\|^{2\nu}+\|A\|^{2\nu})\epsilon_{A}+n^{-\nu}\sup_{j\leq J}{|\sigma_{j}^{2\nu}-\tilde{\sigma}_{j}^{2\nu}|}$
@@
Note that the initial error $x_{\dagger}-x_{0}\in\overline{\mathrm{Range}(A_{\dagger}^{*})}$. The polar decomposition $A_{\dagger}=Q(A_{\dagger}^{*}A_{\dagger})^{\frac{1}{2}}$ with a partial isometry $Q$ (i.e., $Q^{*}Q$ and $QQ^{*}$ are projections) implies $x_{\dagger}-x_{0}\in\overline{\mathrm{Range}\big((A_{\dagger}^{*}A_{\dagger})^{\frac{1}{2}}\big)}$. Thus, for any $\epsilon_{0}>0$, there exists some $\tilde{x}_{0}$, satisfying Assumption 2.1(ii) with $\nu=\frac{1}{2}$ and $w\equiv w_{\epsilon_{0}}\in Y$ (which depends on $\epsilon_{0}$), such that $\|x_{0}-\tilde{x}_{0}\|<\epsilon_{0}$. Let $\tilde{x}^{\delta}_{k}$ be the $k$th (r)SVRG iterate starting with $\tilde{x}_{0}$ and $\tilde{e}^{\delta}_{k}=\tilde{x}^{\delta}_{k}-x_{\dagger}$. Then, for any $b\in(0,1)$, by Lemma A.1 and the inequality (3.5), we can bound $\|P^{k}e_{0}^{\delta}\|$ in (3.3) and (3.8) by
@@
$\leq\big((2c_{0})^{-\frac{1}{2}}k^{-\frac{1}{2}}+n^{-\frac{1}{2}}a\delta^{b}\big)\|w\|+\epsilon_{0}.$
Consequently, for some constant $c^{*}$ depending on $\epsilon_{0}$, there hold
$\begin{aligned} \mathbb{E}[\|e_{k}^{\delta}\|^{2}]^{\frac{1}{2}}&\leq\epsilon_{0}+c^{*}k^{-\frac{1}{2}}+c^{*}\left\{\begin{array}[]{cc}\delta^{\min(b,1-b)},&a>0,\\ n^{-\frac{1}{2}}\sqrt{k}\delta,&a=0,\end{array}\right.\quad c_{0}<\overline{C_{0}},\\ \|e_{k}^{\delta}\|&\leq\epsilon_{0}+\sqrt{n}c^{*}k^{-\frac{1}{2}}\ln k+c^{*}\left\{\begin{array}[]{cc}\delta^{\min(b,1-b)},&a>0,\\ n^{-\frac{1}{2}}\sqrt{k}\delta,&a=0,\end{array}\right.\quad c_{0}<C_{0}.\end{aligned}$
Now for any fixed $\varepsilon>0$, there exists some $\tilde{x}_{0}$ such that $\epsilon_{0}<\frac{1}{3}\varepsilon$. Given the constant $c^{*}$ (depending on $\epsilon_{0}$), there exists some $k_{1}\in\mathbb{N}$ such that for any $k\geq k_{1}$, we have $\sqrt{n}c^{*}k^{-\frac{1}{2}}\ln k<\frac{1}{3}\varepsilon$. Now we discuss the two cases (i) $a>0$ and $b\in(0,1)$, and (ii) $a=0$ separately.
Case (i) $a>0$ and $b\in(0,1)$. Then there exists some $\delta_{1}>0$ such that for any $\delta\leq\delta_{1}$, there holds $c^{*}\delta^{\min(b,1-b)}<\frac{1}{3}\varepsilon$, which implies that for any $k\geq k_{1}$, there hold
$\lim_{\delta\to 0^{+}}\mathbb{E}[\|e_{k}^{\delta}\|^{2}]^{\frac{1}{2}}$ $<\varepsilon,\quad\mbox{if }c_{0}<\overline{C_{0}},$ (3.9)
$\lim_{\delta\to 0^{+}}\|e_{k}^{\delta}\|$ $<\varepsilon,\quad\mbox{if }c_{0}<C_{0}.$ (3.10)
Case (ii) $a=0$. Similar to case (i), when $a=0$, there exists some $\delta_{2}>0$ such that for any $\delta\leq\delta_{2}$, there holds $c^{*}n^{-\frac{1}{2}}\sqrt{k_{1}}\delta<\frac{1}{3}\varepsilon$. Further, for any fixed $\delta\leq\delta_{2}$, there exists some $k_{2}(\delta)\geq k_{1}$ such that for any $k_{1}\leq k(\delta)\leq k_{2}(\delta)$, there hold
$\sqrt{n}c^{*}k(\delta)^{-\frac{1}{2}}\ln k(\delta)<\tfrac{1}{3}\varepsilon\quad\mbox{and}\quad c^{*}n^{-\frac{1}{2}}\sqrt{k(\delta)}\delta<\tfrac{1}{3}\varepsilon.$
This discussion implies that for the stopping index $k(\delta)$ satisfying the following conditions
$\lim_{\delta\to 0^{+}}k(\delta)=\infty\quad\mbox{and}\quad\lim_{\delta\to 0^{+}}\sqrt{k(\delta)}\delta=0,$
there again hold the relations (3.9) and (3.10).
Combining the preceding discussions completes the proof of the corollary. ∎